# Backtracking New Q-Newton’s method for finding roots of meromorphic functions in 1 complex variable: Global convergence, and local stable/unstable curves

### v1

@@ Abstract.
If $f=P/Q$, where $P$ and $Q$ are polynomials in 1 complex variable z with $\deg(P)>\deg(Q)$, we show the existence of an exceptional set $\mathcal{E}\subset\mathbf{C}$, which is contained in a countable union of real analytic curves in $\mathbf{R}^{2}=\mathbf{C}$, so that the following statements A and B hold. Here, $\{z_{n}\}$ is the sequence constructed by BNQN with an initial point $z_{0}$, not a pole of $f$.
[…]
@@
Based on the theoretical results, we explain (both rigorously and heuristically) of what observed in experiments with BNQN, in previous works by the authors. The dynamics of BNQN seems also to have some similarities (and differences) to the classical Leau-Fatou’s flowers.
@@ 1. Introduction
where $\gamma\in\mathbf{C}$ is a small enough constant, has better global convergence guarantee for polynomials. In fact, [11] and [18] show that if certain conditions are satisfied for the polynomial $f$ (or more generally a rational function), then global convergence to roots is achieved. The idea is to relate this to Newton’s flow method. (For a good survey on Relaxed Newton’s method, Newton’s flow and some other methods, the reader can consult [4].) On the other hand, in practice can have less than optimal rate of convergence.
@@
[…]
@@
Experiments, for simple polynomials of degree $4$, seem to indicate that in general $\mathcal{E}$ cannot be contained in a finite union of real analytic curves, see Figures 8 and 9. On the other hand, pictures seem to indicate that $\mathcal{E}$ should be locally contained in a finite union of real analytic curves. We note that the proof of Theorem 1.2 constructs explicitly certain subsets of $\mathcal{E}$ which are contained in a finite number of real analytic curves and whose union is $\mathcal{E}$. Moreover, we have the following characterisation.
@@
The next picture, Figure 1, compares the performance of Newton’s method and BNQN for finding the 7 roots, within the domain $|z|\leq 10$, of $f(z)=e^{2iz}-1$, an entire function without critical points. We also draw a picture for Voronoi’s diagram of the corresponding 7 roots.
Figure 1. Basins of attraction for finding the 7 roots inside the domain $|z|<10$ of $f(z)=e^{2iz}-1$, an entire function with no critical points. Top picture: Voronoi’s diagram of the 7 roots, center picture: Newton’s method, bottom picture: BNQN. Points of the same colour belongs to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 7 roots. Here $\theta=1$ in Algorithm 1.
The second main result of this paper is the following, which is the main tool to prove Theorem 1.2. For $r>0$, we use the usual notation $\mathbb{D}(0,r)=\{z\in\mathbf{C}:|z|<r\}$. We refer to Section 2 for some details on stable and unstable manifolds/curves of dynamical systems. In the statement of the theorem, $arg(z)$ is the branch of the argument function with values in $(0,2\pi]$.
@@
Then there is $r_{0}>0$, and continuous curves $C_{1},\ldots,C_{2d}$ in $\mathbb{D}(0,r_{0})$, with the following properties:
@@
2) The domains $U_{j}$’s ($j=1,\ldots,d$) in Theorem 1.4 are reminiscent of the classical Leau-Fatou’s flowers (please see [19] for this notion, see also the original works in [15], [13], [8]) for 1-dimensional dynamics near a parabolic fixed point of a holomorphic germ $p(z)=z+az^{d+1}+$ higher order terms, where $0\not=a\in\mathbf{C}$ is a constant. Section 3 contains a detailed discussion on this.
@@
Similarly, it can be shown that when $f$ is a polynomial, BNQN applied to $F(x,y)=|f(x+iy)|^{2}$ is almost a linear map near $\infty$. This is similar to the usual Newton’s method.
We also remark that as a byproduct of the work done in this paper, specifically the proof of Lemma 3.2, since $z_{k}-v_{k}$ in that lemma is (up to an $O(|z_{k}|^{2})$ error) the update rule for Newton’s method for Optimization (for detail please see e.g. [9]) and it is equal to
@@
it is rigorously seen that for Newton’s method for Optimization applied to $F(x,y)=|f(x+iy)|^{2}$, the critical points (no matter if it is a root or not) of $f(z)$ are also attractors, as observed experimentally in [9]. Therefore, like Newton’s method, Newton’s method for Optimization also does not have global convergence guarantee for finding roots of polynomials in 1-complex variable.
The plan for the remaining of this paper is as follows. In Section 2 we present some preliminary results. In Section 3 we prove Theorem 1.4, and in Section 4 we prove Theorems 1.2, Theorem 1.3 and 1.6. Also, in Section 3 we discuss the relation between the dynamics of BNQN near a critical point and the classical Leau-Fatou’s flowers of holomorphic dynamics near a parabolic fixed point, and in Section 4 we discuss how the results in this paper can be used to explain (both rigorously and heuristically) the observations in many experiments on BNQN, in particular those in [9].
@@ 2.1. BNQN and some first properties
Given: $\{\delta_{0},\delta_{1},\ldots,\delta_{m}\}\subset\mathbf{R}$,  $\tau\in 2\mathbf{N}$, and $0<\gamma_{0}\leq 1$;
@@ 2.2. Some results on stable/unstable manifolds of real analytic maps in dimension 2
For the first result, there are proofs for invertible germs, but we cannot find a reference where it is proved for non-invertible germs. We will provide a sketch of proof for it, sufficiently self-contained, following those in [24, Chapters 5 and 6] (differentiable case) and [20, Section 6.4] (complex analytic case).
@@
First, complexify the real variables $x,y$ to complex variables $X,Y$, and shrink the domain if necessary, we obtain a complex analytic map $\widehat{H}:\widehat{U}\rightarrow\mathbf{C}^{2}$. The Jacobian of $\widehat{H}$ at $(0,0)$ has the same eigenvalues and eigenvectors as $H$. By [20, Theorem 6.4.3, case ii], $\widehat{H}$ has an unstable manifold $\widehat{C}$, which is given by an equation of the form $X=\varphi(Y)$. Take the intersection of this graph with $\mathbf{R}^{2}$ we obtain the unstable manifold $C_{2}$ for $H$, which is real analytic. Moreover, we can, after a linear change of coordinates, assume that the unstable manifold of $\widehat{H}$ is the $Y$ axis, and $\widehat{H}(0,Y)=(0,\lambda_{2}Y)$.
@@
Assume $\Omega\subset\mathbf{R}^{2}$ is a small open neighbourhood of 00, and $h:\Omega\rightarrow\mathbf{R}^{2}$ is a real analytic map. Assume that $Jh(0,0)$ has two eigenvalues $a,A$ such that $0\leq|a|<1<|A|$.
@@
Since the 2 eignevalues $a,A$ of $Jh(0,0)$ are different, by applying a linear change of coordinates, we can assume that the following holds. $Jh(0,0)$ is a diagonal matrix with stable eigenvector $(1,0)$, stable eigenvalue $10<a<$ 1, and unstable eigenvector $(0,1)$, unstable eigenvalue $A>1$. Using Theorem 2.3, we can also assume that the stable manifold is of the form $y=u(x)$ for a real analytic function $u$ with $u(0)=u^{\prime}(0)=0$, and that the unstable manifold is of the form $x=v(y)$ for a real analytic function $v$ with $v(0)=v^{\prime}(0)=0$. In this case, the stable and unstable manifolds partition $\Omega$ into the following quadrants divided by the stable and unstable manifolds (see Figure 2):
@@
In this section we research the local behaviour of BNQN near a critical point $z^{*}=0$ but not a root of $f$, which accumulates in the proof of Theorem 1.4. We also point out some similarities and differences between the dynamics of BNQN near $z^{*}$ and the classical Leau-Fatou’s flowers of a holomorphic germ in 1 complex variable near a parabolic fixed point.
@@
We can, either by using Theorem 2.2 or by mimicking the calculations to be done in this section, assume throughout this section that $f(z)$ has the form:
@@ 3.1. Some local calculations
$U_{1}=\left[\begin{array}[]{ccc}\cos((d-2)\theta/2)+O(r)\\ -\sin((d-1)\theta/2)+O(r)\\ \end{array}\right],$
@@ 3.3. The dynamics of the linearization $\Phi_{1}(z)$
Figure 4. The angle in $U_{1}$, the sector bound by the stable ray $C_{2d}$ and the unstable ray $C_{1}$, is increasing under $\Phi_{1}$.
@@ 3.4.1. The case $d\geq 3$
Figure 5. Transforming the small sector into a half of a rectangle containing $(0,0)$ inside. The curves $C_{2d}^{1}$, $C_{2d}^{2}$ in the sector are pullback of certain curves $Z_{0_{1}}$, $Z_{0_{2}}$ in the rectangle. To make sure that $C_{2d}^{2}$ does not cross the boundary of the sector, we need to shrink both the sector and the rectangle.
@@
with the corresponding eigenvectors $(1,0)$ and $(0,1)$. Therefore, by shrinking $r_{0}$ and $\varepsilon_{1}$ (also $\varepsilon_{0}$ so that still we have $\psi(S_{r_{0},\varepsilon_{0},2d})\subset R_{r_{0},\varepsilon_{1}}$), we can assume that $(0,0)$ is the only fixed point of $H$ inside $R_{r_{0},\varepsilon_{1}}$ and both Theorems 2.3 and 2.4 apply. (In this case, instead of Theorem 2.3, we can do as follows. First, complexify and use [20, Theorem 6.4.1] to obtain complex stable and unstable manifolds, then use [24] to deduce that $C^{1}$ stable and unstable manifolds for the original map exist, therefore the intersection of the complex stable and unstable manifolds to the real plane are smooth real analytic curves.)
@@
This has eigenvalues $0=\lambda_{1}<1<\frac{d}{d-1}=\lambda_{2}$, with eigenvectors $(1,0)$ and $(0,1)$. hence Theorem 2.3 provides an unstable manifold which is real analytic. (In this case, instead of Theorem 2.3, we can do as follows. First, complexify and use [20, Theorem 6.4.3 case ii] to obtain the complex unstable manifold, then use [24] to deduce that $C^{1}$ unstable manifold for the original map exists, therefore the intersection of the unstable manifold to the real plane is a smooth real analytic curve.) Again, as in Step 3, the pullback of the unstable curve of the map $\widehat{H}$ can be extended to a closed smooth real analytic curve in $\mathbb{D}(0,r_{0})$ which is tangent to the x-axis at $(0,0)$. Pullback further by the rotation $\varphi$, we obtain the real analytic curves $\widehat{C_{1}}$ and $C_{1}$ in Theorem 1.4.
@@ 3.4.2. Proof for the case $d=2$
3.5. Some similarities and differences between BNQN and Leau-Fatou’s flowers
We now point out some remarkable similarities, and differences, between the behaviour of BNQN near a critical point but not a root of $f(z)$ and the classical Leau-Fatou’s flowers. The interested reader can consult more in [19], and also the original works in [15], [13], [8]. One considers a holomorphic germ $p(z)=z+az^{d+1}+$ higher order terms. This is a parabolic fixed point. The holomorphic dynamics of $p(z)$ near 00, and the dynamics near 00 of BNQN applied to $F(z)=|f(z)|^{2}$ where $f(z)=1+az^{d}+$ higher order terms - have some remarkable similarities. However, there are also clear differences. We will discuss both similarities and differences in the below, for $d\geq 2$.
Similarities:
1) $p(z)=zf(z)$.
2) In both cases, there are $d$ stable vectors and $d$ unstable vectors. The angle between a vector of a type and one of its neighbours of the opposite type is $\pi/d$.
3) There are ”repelling petals”, each being a domain bound by two consecutive stable curves and contains a unique unstable curve inside. Similarly, there are also ”attracting petals”, each being a domain bound by two consecutive unstable curves and contains a unique stable curve inside..
Differences:
1) A parabolic fixed point is partially attracting and partially repelling for the holomorphic dynamics of $p(z)$, while a critical point (but not root) of $f$ is a (degenerate) saddle point for BNQN.
2) ”Attracting petals” are invariant and attracting for the holomorphic dynamics $p(z)$. ”Repelling petals” are invariant and repelling for the BNQN dynamics.
In the literature, the holomorphic dynamics $p(z)$ on the ”repelling petals” is not extensively explored. ”Attracting petals” are invariant but still repelling for the BNQN dynamics.
3) The union of attracting petals (boundaries included), in the Leau-Fatou’s flowers picture, is not an open neighbourhood of 00. The union of repelling petals (with boundaries included), for BNQN, is an open neighbourhood of 00.
4. Proofs of Theorems 1.2, Theorem 1.3 and 1.6
We provide in this section proofs of Theorems 1.2, Theorem 1.3 and 1.6, and a couple of experiments indicating that the condition that $f$ has compact sublevels is needed for the conclusions of Part II of Theorem 1.2. We finish the section with a discussion on how the results in this paper can be used to heuristically explain the apparent connections between BNQN, Voronoi’s diagram and Newton’s flow, as well as the boundary of basins of attraction for roots in BNQN dynamics seems to be smooth, in particular from many experiments done in [10].
@@ 3.5. Some similarities and differences between BNQN and Leau-Fatou’s flowers
Curves of type 1: the boundaries of the discs, or the boundaries of the small sectors in these discs. By the proof of Theorems 1.4 and 1.6, inside each of these discs, BNQN is real analytic outside curves of type 1. Since $\Omega$ is bounded, only a finite number of such points $z^{*}$, and hence of curves of type 1, which is relevant.
@@
Now, let $S$ be the union of the curves of type 1, 2, 3, and 4, we observe that for $z_{0}\in K\backslash S$ there is no ”switch” in the dynamics of BNQN: if at $z_{0}$ the parameters in Algorithm 1 are $\delta_{j}$ and $\gamma$, then in an open neighbourhood of $z_{0}$, BNQN still uses the same parameters. As in the proof of Theorem 1.4, we can show that when there is no switch then BNQN is real analytic, since the involved matrix $A(z)$ in Algorithm 1 is real analytic and invertible.

### v2

@@ Abstract.
If $f=P/Q$, where $P$ and $Q$ are polynomials in 1 complex variable z with $\deg(P)>\deg(Q)$, we show the existence of an exceptional set $\mathcal{E}\subset\mathbf{C}$, which is contained in a countable union of real analytic curves in $\mathbf{R}^{2}=\mathbf{C}$, so that the following statements A and B hold. Here, $\{z_{n}\}$ is the sequence constructed by BNQN with an initial point $z_{0}$ which is not a pole of $f$.
A) If $\𝐂z_{0}\in\mathbf{C}\backslash\mathcal{E}$, then $\{z_{n}\}$ converges to a root of $f$.
@@
Based on the theoretical results, we explain (both rigorously and heuristically) of what observed in experiments with BNQN, in previous works by the authors. In particular, the dynamics of BNQN (an iterative method) seems to have some striking similarities to Newton’s method (a continuous method) and the classical Poincaré-Bendixon theorem for differentiable real dynamical systems on the complex plane. This is the more interesting given that discrete versions of Newton’s method (e.g. Relaxed Newton’s method) does not behave this way.
@@ 1. Introduction
where $\gamma\in\mathbf{C}$ is a small enough constant, has better global convergence guarantee for polynomials. In fact, [12] and [17] show that if certain conditions are satisfied for the polynomial $f$ (or more generally a rational function), then global convergence to roots is achieved. The idea is to relate this to Newton’s flow method. (For a good survey on Relaxed Newton’s method, Newton’s flow and some other methods, the reader can consult [5].) On the other hand, in practice can have less than optimal rate of convergence. Indeed, there seems to be a striking similarity between the behaviour of BNQN (an iterative method) and that of Newton’s method (a continuous method) and the classical Poincaré-Bendixon theorem [21][4], which will be discussed in more detail in Section 4.4.
@@
2) A set $A\subset\mathbf{C}$ is a real analytic curve if there is an open subset $U\subset\mathbf{C}$ so that $A$ is defined by a real analytic equation in $U$, and has dimension $1$.
@@
Experiments, for simple polynomials of degree $4$, seem to indicate that in general $\mathcal{E}$ cannot be contained in a finite union of real analytic curves, see Figures 10 and 11. On the other hand, pictures seem to indicate that $\mathcal{E}$ should be locally contained in a finite union of real analytic curves. We note that the proof of Theorem 1.2 constructs explicitly certain subsets of $\mathcal{E}$ which are contained in a finite number of real analytic curves and whose union is $\mathcal{E}$. Moreover, we have the following characterisation.
@@
Figures 1 and 2 compare the performance of Newton’s method, Random Relaxed Newton’s method and BNQN for finding the 7 roots, within the domain $|z|\leq 10$, of $f(z)=e^{2iz}-1$, an entire function without critical points. We also draw a picture for Voronoi’s diagram of the corresponding 7 roots.
Figure 1. Basins of attraction for finding the 7 roots inside the domain $|z|<10$ of $f(z)=e^{2iz}-1$, an entire function with no critical points. Top picture: Newton’s method, bottom picture: Random Relaxed Newton’s methods. Points of the same colour belong to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 7 roots. Note that the 2 pictures look very similar.
Figure 2. Basins of attraction for finding the 7 roots inside the domain $|z|<10$ of $f(z)=e^{2iz}-1$, an entire function with no critical points. Top picture: Voronoi’s diagram of the 7 roots, bottom picture: BNQN. Points of the same colour belong to basin of attraction of the same root. Points in black do not belong to the basin of attraction of any of the concerned 7 roots. Here $\theta=1$ in Algorithm 1. Note that the 2 pictures look very similar.
The second main result of this paper is the following, which is the main tool to prove Theorem 1.2. For $r>0$, we use the usual notation $\mathbb{D}(0,r)=\{z\in\mathbf{C}:|z|<r\}$. We refer to Section 2 for some details on stable and unstable manifolds/curves of dynamical systems. In the statement of the theorem, $arg(z)$ is the branch of the argument function with values in $(0,2\pi]$. See Figure 3 for an illustration of Theorem 1.4 when $d=3$.
Figure 3. An illustration of Theorem 1.4: The orbits of BNQN near a critical point of order $d=3$. Here the local dynamics is that of a 3-fold saddle point.
@@
The point $z^{*}$ is a d-fold saddle point for the dynamics of BNQN. More precisely, there are $r_{0}>0$, and continuous curves $C_{1},\ldots,C_{2d}$ in $\mathbb{D}(0,r_{0})$, with the following properties:
@@
2) The behaviour of BNQN described in Theorem 1.4 is very similar to that of Newton’s flow. The latter can be explained through the classical Poincaré-Bendixon theorem [21][4]. We will discuss more on this in Section 4.4.
@@
As a byproduct of the proof of Lemma 3.2, since $z_{k}-v_{k}$ in that lemma is (up to an $O(|z_{k}|^{2})$ error) the update rule for Newton’s method for Optimization (for detail please see e.g. [10]) and it is equal to
@@
it is rigorously seen here that for Newton’s method for Optimization applied to $F(x,y)=|f(x+iy)|^{2}$, the critical points (no matter if it is a root or not) of $f(z)$ are also attractors, as observed experimentally in [10]. Therefore, like Newton’s method, Newton’s method for Optimization also does not have global convergence guarantee for finding roots of polynomials in 1-complex variable.
Plan of the paper. In Section 2 we present some preliminary results. In Section 3 we prove Theorem 1.4, and in Section 4 we prove Theorems 1.2, Theorem 1.3 and 1.6. Also, in Section 4 we discuss how the results in this paper can be used to explain (both rigorously and heuristically) the observations in many experiments on BNQN, in particular those in [10], in particularly connections to Newton’s flow and the classical Poincaré-Bendixon theorem.
@@ 2.1. BNQN and some first properties
Given: $\{\delta_{0},\delta_{1},\ldots,\delta_{m}\}\subset\mathbf{R}$,  $\tau\in 2\mathbf{N}$, $0\leq\theta$, and $0<\gamma_{0}\leq 1$;
@@ 2.2. Some results on stable/unstable manifolds of real analytic maps in dimension 2
For the first result, there are proofs for invertible germs, but we cannot find a reference where it is proved for non-invertible germs. We will provide a sketch of proof for it, sufficiently self-contained, following those in [24, Chapters 5 and 6] (differentiable case) and [19, Section 6.4] (complex analytic case).
@@
First, complexify the real variables $x,y$ to complex variables $X,Y$, and shrink the domain if necessary, we obtain a complex analytic map $\widehat{H}:\widehat{U}\rightarrow\mathbf{C}^{2}$. The Jacobian of $\widehat{H}$ at $(0,0)$ has the same eigenvalues and eigenvectors as $H$. By [19, Theorem 6.4.3, case ii], $\widehat{H}$ has an unstable manifold $\widehat{C}$, which is given by an equation of the form $X=\varphi(Y)$. Take the intersection of this graph with $\mathbf{R}^{2}$ we obtain the unstable manifold $C_{2}$ for $H$, which is real analytic. Moreover, we can, after a linear change of coordinates, assume that the unstable manifold of $\widehat{H}$ is the $Y$ axis, and $\widehat{H}(0,Y)=(0,\lambda_{2}Y)$.
@@
Assume $\Omega\subset\mathbf{R}^{2}$ is a small open neighbourhood of 00, and $h:\Omega\rightarrow\mathbf{R}^{2}$ is a real analytic map. Assume that $Jh(0,0)$ has two eigenvalues $a,A\in\mathbf{R}$ such that $0\leq|a|<1<|A|$.
@@
Since the 2 eigenvalues $a,A$ of $Jh(0,0)$ are different, by applying a linear change of coordinates, we can assume that the following holds. $Jh(0,0)$ is a diagonal matrix with stable eigenvector $(1,0)$, stable eigenvalue $10<a<$ 1, and unstable eigenvector $(0,1)$, unstable eigenvalue $A>1$. Using Theorem 2.3, we can also assume that the stable manifold is of the form $y=u(x)$ for a real analytic function $u$ with $u(0)=u^{\prime}(0)=0$, and that the unstable manifold is of the form $x=v(y)$ for a real analytic function $v$ with $v(0)=v^{\prime}(0)=0$. In this case, the stable and unstable manifolds partition $\Omega$ into the following quadrants divided by the stable and unstable manifolds (see Figure 4):
@@
In this section we research the local behaviour of BNQN near a critical point $z^{*}=0$ but not a root of $f$, which accumulates in the proof of Theorem 1.4.
@@
We can, either by using Theorem 2.2 or by mimicking the arguments in this section, assume throughout this section that $f(z)$ has the form:
@@ 3.1. Some local calculations
$U_{1}=\left[\begin{array}[]{ccc}\cos((d-2)\theta/2)+O(r)\\ -\sin((d-2)\theta/2)+O(r)\\ \end{array}\right],$
@@ 3.3. The dynamics of the linearization $\Phi_{1}(z)$
Figure 6. The angle in $U_{1}$, the sector bound by the stable ray $C_{2d}$ and the unstable ray $C_{1}$, is expanding under $\Phi_{1}$.
@@ 3.4.1. The case $d\geq 3$
Figure 7. Transforming the small sector into a half of a rectangle containing $(0,0)$ inside. The curves $C_{2d}^{1}$, $C_{2d}^{2}$ in the sector are pullback of certain curves $Z_{0_{1}}$, $Z_{0_{2}}$ in the rectangle. To make sure that $C_{2d}^{2}$ only exits the sector through the boundary on the right side (i.e. an arc of the circle), we need to shrink both the sector and the rectangle (from the big red rectangle to the smaller blue rectangle).
@@
with the corresponding eigenvectors $(1,0)$ and $(0,1)$. Therefore, by shrinking $r_{0}$ and $\varepsilon_{1}$ (also $\varepsilon_{0}$ so that still we have $\psi(S_{r_{0},\varepsilon_{0},2d})\subset R_{r_{0},\varepsilon_{1}}$), we can assume that $(0,0)$ is the only fixed point of $H$ inside $R_{r_{0},\varepsilon_{1}}$ and both Theorems 2.3 and 2.4 apply. (In this case, instead of Theorem 2.3, we can do as follows. First, complexify and use [19, Theorem 6.4.1] to obtain complex stable and unstable manifolds, then use [24] to deduce that $C^{1}$ stable and unstable manifolds for the original map exist, therefore the intersection of the complex stable and unstable manifolds to the real plane are smooth real analytic curves.)
@@
This has eigenvalues $0=\lambda_{1}<1<\frac{d}{d-1}=\lambda_{2}$, with eigenvectors $(1,0)$ and $(0,1)$. hence Theorem 2.3 provides an unstable manifold which is real analytic. (In this case, instead of Theorem 2.3, we can do as follows. First, complexify and use [19, Theorem 6.4.3 case ii] to obtain the complex unstable manifold, then use [24] to deduce that $C^{1}$ unstable manifold for the original map exists, therefore the intersection of the unstable manifold to the real plane is a smooth real analytic curve.) Again, as in Step 3, the pullback of the unstable curve of the map $\widehat{H}$ can be extended to a closed smooth real analytic curve in $\mathbb{D}(0,r_{0})$ which is tangent to the x-axis at $(0,0)$. Pullback further by the rotation $\varphi$, we obtain the real analytic curves $\widehat{C_{1}}$ and $C_{1}$ in Theorem 1.4.
@@ 3.4.2. Proof for the case $d=2$
4. Proofs of Theorems 1.2, 1.3 and 1.6
We provide in this section proofs of Theorems 1.2, 1.3 and 1.6. We finish the section, with some illustrating experiments and a discussion on how the results in this paper can be used to heuristically explain the apparent connections between BNQN, Voronoi’s diagram and Newton’s flow, as well as the boundary of basins of attraction for roots in BNQN dynamics seems to be smooth, in particular from many experiments done in [11], in particularly to Newton’s flow and Poincaré-Bendixon theorem.
@@
Curves of type 0: the boundaries of the discs, or the boundaries of the small sectors in these discs. By the proof of Theorems 1.4 and 1.6, inside each of these discs, BNQN is real analytic outside curves of type 1. Since $\Omega$ is bounded, only a finite number of such points $z^{*}$, and hence of curves of type 1, which is relevant.
@@
Now, let $S$ be the union of the curves of type 0, 1, 2, 3, and 4, we observe that for $z_{0}\in K\backslash S$ there is no ”switch” in the dynamics of BNQN: if at $z_{0}$ the parameters in Algorithm 1 are $\delta_{j}$ and $\gamma$, then in an open neighbourhood of $z_{0}$, BNQN still uses the same parameters. As in the proof of Theorem 1.4, we can show that when there is no switch then BNQN is real analytic, since the involved matrix $A(z)$ in Algorithm 1 is real analytic and invertible.